OpenStudy (anonymous):

1.97 + _______ = $5.00*

OpenStudy (anonymous):

3.03

Nnesha (nnesha):

1.97 - 5 = 3.03 so 1.97 + 3.03 = 5

OpenStudy (anonymous):

5.00 - 1.97 = your answer